# relsizeFont size

### v2

@@ B.1 Supervised Learning
B.2 Active learning
We use the same setting as training on ImageNet-1K in Sec. LABEL:sub@subsec:appendix_supervised.
B.3 Continual learning
@@
Appendix C Selection Hyperparameter Settings
C.1 Selection Hyperparameters
We conducted a grid search to optimize three hyperparameters: the number of neighbors $k$ for constructing the $k$NN graph, the cutoff ratio $\beta$ to remove the most difficult samples, and the imbalance factor $\gamma$ to maintain the balance between different classes.
@@ C.2 Preliminary Experiments on k
C.3 Grid Search Results
For all experiments, we search $\beta$ in $\{-1,-0.95,\dots,-0.05,0,0.05,\dots,0.95,1\}$ and $\gamma$ in $\{1,1.05,\dots,1.5\}$. For CIFAR10 and CIFAR100, we search $k$ in $\{10,11,\dots,30\}$. For ImageNet-1K, we search it only in $\{20,25\}$ due to computational costs. To save computational costs, we set $k$ to around $\log_{2}n$ for experiments on other datasets. We evaluate every possible combination of the hyperparameters. Table LABEL:sub@tab:hyper presents the optimal hyperparameters for supervised learning and active learning. Table LABEL:sub@tab:continual_hyper presents the optimal hyperparameters for continual learning.
@@ C.3 Grid Search Results
Appendix D Detailed Experimental Results
We present detailed experimental results for supervised learning from Table LABEL:sub@tab:cls to LABEL:sub@tab:vqa. Our method consistently performs better than baselines across all tasks and sampling rates. Meanwhile, several factors prevent us from achieving more significant improvements over the baselines in certain settings:
  • In high-rate settings for simple datasets, such as the $70\%$ setting for CIFAR10, all methods perform closely to using the entire training set, leaving little room for improvement.
  • In low-rate settings for challenging datasets, such as the $1\%$ setting for ImageNet-1K, the performance of all methods is limited due to the insufficient number of samples per class.
  • In scenarios where models are fine-tuned from pretrained weights, the knowledge within in these models anchors the performance at a certain level, leading to similar performance across all methods.
@@ Appendix D Detailed Experimental Results
Table 10: Results on image classification. The best one is bold, and the runner-up is underlined.
Dataset CIFAR10 (100%:95.49) CIFAR100 (100%:77.90) ImageNet-1K (100%:73.63)
Sampling rate 70% 50% 20% 10% 5% 2% 1% 70% 50% 20% 10% 5% 2% 1% 70% 50% 20% 10% 5% 2% 1%
Random 94.29 92.33 84.17 71.95 59.88 45.51 35.44 74.72 70.51 52.29 37.38 23.03 13.38 8.57 71.63 69.26 58.90 47.10 34.04 16.56 5.50
Moderate 94.04 92.36 80.96 63.91 49.94 32.00 26.18 74.61 69.95 49.37 30.53 17.68 9.28 5.97 71.33 68.72 55.23 40.97 25.75 11.33 4.52
CCS 94.23 93.83 83.85 75.56 65.89 49.02 40.24 76.10 73.12 58.39 44.92 28.68 15.93 10.28 70.74 69.23 60.04 50.41 36.92 19.92 9.43
$\mathbb{D}^{2}$ Pruning 93.17 93.70 86.83 76.56 65.12 45.44 39.03 76.00 74.45 59.08 44.78 26.68 13.90 9.61 71.29 70.32 58.91 50.81 37.12 18.97 11.23
GraphCut 94.38 92.43 81.98 69.29 59.36 46.93 39.68 76.64 68.79 25.08 16.46 10.78 7.84 5.94 68.91 68.72 55.28 44.79 33.54 20.07 11.49
Entropy 94.13 92.39 75.34 60.66 46.17 35.85 29.28 76.81 62.52 33.28 21.85 13.90 8.31 5.30 70.93 69.21 54.76 38.46 22.78 7.01 1.95
Forgetting 94.76 94.34 57.72 35.21 30.87 27.22 23.62 76.64 68.75 25.17 16.45 10.70 7.82 5.90 70.57 70.46 60.77 48.73 33.86 15.13 5.66
EL2N 94.75 94.10 46.41 22.40 15.66 13.05 12.75 76.15 66.19 14.85 7.50 5.28 3.83 3.17 71.68 65.98 31.90 12.57 6.50 3.25 1.90
AUM 94.90 94.38 51.98 30.56 22.92 18.10 14.50 76.84 67.81 16.66 8.54 5.60 4.35 3.56 69.94 65.36 21.91 10.50 6.42 3.58 2.24
Variance 90.45 85.81 64.90 53.64 44.45 36.61 31.64 73.69 68.42 49.60 34.63 22.70 13.71 9.18 70.12 66.09 35.15 13.85 7.13 4.72 1.81
SES (Ours) 95.01 94.50 88.31 80.24 69.82 54.78 45.25 77.23 74.63 61.52 48.02 33.39 19.68 14.16 72.80 71.05 63.24 53.59 41.88 25.59 13.43
@@
Table 11: Results on text classification. The best one is bold, and the runner-up is underlined.
Dataset ANLI (100%:49.25) IMDB Review(100%:95.90)
Sampling rate 70% 50% 20% 10% 5% 2% 1% 70% 50% 20% 10% 5% 2% 1%
Random 47.08 45.20 42.13 39.52 38.82 37.50 35.96 95.25 95.04 93.53 91.89 89.60 83.25 73.31
Moderate 46.84 45.11 41.95 40.16 38.99 35.83 33.91 95.39 95.31 94.03 92.55 90.34 58.34 50.81
CCS 46.56 45.92 41.67 41.63 40.33 37.41 36.82 95.42 95.35 93.44 91.87 89.89 85.73 82.05
$\mathbb{D}^{2}$ Pruning 48.56 47.49 42.77 41.43 40.34 37.92 36.29 95.43 95.40 93.75 92.44 90.77 87.40 80.06
GraphCut 46.14 44.53 42.12 39.86 38.15 35.44 34.02 95.39 95.21 93.37 91.69 90.93 86.85 82.26
Entropy 46.32 45.53 41.45 39.67 38.54 36.69 36.40 95.39 95.17 93.92 92.57 90.30 77.16 59.41
Forgetting 48.73 42.29 39.82 38.37 35.95 35.78 35.03 95.04 95.31 93.45 89.25 53.73 57.57 51.92
EL2N 48.70 47.85 43.14 39.63 37.52 34.33 34.27 95.29 95.34 91.31 60.29 49.88 47.18 43.74
AUM 47.86 47.58 43.57 40.02 34.66 34.16 33.62 95.27 95.23 90.60 55.68 49.81 43.13 36.29
Variance 47.97 47.87 40.70 38.75 33.52 33.50 33.17 95.44 95.40 92.34 49.76 48.95 44.90 45.53
SES (Ours) 49.00 48.22 45.94 43.63 41.82 39.88 38.16 95.60 95.42 94.54 92.96 91.42 88.58 84.04
@@
Random 74.02 72.10 65.45 57.56 43.47 18.78 9.24
Moderate 73.42 72.03 65.12 54.71 40.20 15.97 5.13
CCS 74.64 72.27 65.72 57.35 39.01 17.26 8.49
$\mathbb{D}^{2}$ Pruning 74.46 72.55 65.59 55.73 44.04 19.16 10.75
GraphCut 67.45 64.15 53.12 38.29 26.81 8.56 8.16
AL-MDN 74.51 70.36 65.26 54.51 30.85 12.33 8.97
SES (Ours) 75.20 73.33 66.52 59.52 45.92 23.39 16.15

### v3

@@ D.1 Supervised Learning
D.2 Active learning
We use the same setting as training on ImageNet-1K in Appendix LABEL:sub@subsec:appendix_supervised.
D.3 Continual learning
@@
Appendix E Selection Hyperparameter Settings
E.1 Selection Hyperparameters
We conduct a grid search to optimize three hyperparameters: the number of neighbors $k$ for constructing the $k$NN graph, the cutoff ratio $\beta$ to remove the most difficult samples, and the imbalance factor $\gamma$ to maintain the balance between different classes.
@@ E.2 Preliminary Experiments on k
E.3 Grid Search Results
For all experiments, we search $\beta$ in $\{-1,-0.95,\dots,-0.05,0,0.05,\dots,0.95,1\}$ and $\gamma$ in $\{1,1.05,\dots,1.5\}$. For CIFAR10 and CIFAR100, we search $k$ in $\{10,11,\dots,30\}$. For ImageNet-1K, we search it only in $\{20,25\}$ due to computational costs. To save computational costs, we set $k$ to around $\log_{2}n$ for experiments on other datasets. With this choice, the $k$NN graph has $O(n\log_{2}n)$ edges, resulting in a Shapley value calculation time complexity of $O(n\log_{2}n)$. This complexity is comparable to sorting and is lower than or equal to that of typical sample selection methods, which is at least $O(n\log_{2}n)$ when sorting is required and $O(n^{2})$ when an enumeration of all pairs of samples is required. For extremely large datasets (e.g. DataComp (Gadre et al., 2024) with billions of samples) where this complexity becomes impractical, a potential solution is to split the dataset into bins before selection. Determining an optimal splitting method remains an open question and a promising direction for future research. We evaluate every possible combination of the hyperparameters. Table LABEL:sub@tab:hyper presents the optimal hyperparameters for supervised learning and active learning. Table LABEL:sub@tab:continual_hyper presents the optimal hyperparameters for continual learning.
@@ E.3 Grid Search Results
Appendix F Detailed Experimental Results
F.1 Superivised Learning
We present detailed experimental results for supervised learning from Tables LABEL:sub@tab:app_c10 to LABEL:sub@tab:vqa. Our method consistently performs better than baselines across all tasks and sampling rates. Meanwhile, several factors prevent us from achieving more significant improvements over the baselines in certain settings:
  • In high-sampling-rate settings for simple datasets, such as the $70\%$ setting for CIFAR10, all methods perform closely to using the entire training set, leaving little room for improvement.
  • In low-sampling-rate settings for challenging datasets, such as the $1\%$ setting for ImageNet-1K, the performance of all methods is limited due to the insufficient number of samples per class.
  • In scenarios where models are fine-tuned from pretrained weights, the knowledge within these models anchors the performance at a certain level, leading to similar performance across all methods.
@@ F.1 Superivised Learning
Table 10: Results on image classification (CIFAR10). The best one is bold, and the runner-up is underlined.
Dataset CIFAR10 (100%:95.49)
Sampling rate 70% 50% 20% 10% 5% 2% 1%
Random $94.29\pm 0.08$ $92.33\pm 0.15$ $84.17\pm 0.62$ $71.95\pm 1.50$ $59.88\pm 2.45$ $45.51\pm 1.68$ $35.44\pm 1.49$
Moderate $94.04\pm 0.26$ $92.36\pm 0.17$ $80.96\pm 2.56$ $63.91\pm 1.56$ $49.94\pm 0.70$ $32.00\pm 2.11$ $26.18\pm 1.94$
CCS $94.23\pm 0.35$ $93.83\pm 0.15$ $83.85\pm 1.10$ $75.56\pm 2.39$ $\underline{65.89\pm 1.04}$ $\underline{49.02\pm 1.35}$ $\underline{40.24\pm 0.96}$
$\mathbb{D}^{2}$ Pruning $93.17\pm 0.11$ $93.70\pm 0.20$ $\underline{86.83\pm 0.48}$ $\underline{76.56\pm 1.14}$ $65.12\pm 0.88$ $45.44\pm 1.75$ $39.03\pm 1.69$
GraphCut $94.38\pm 0.20$ $92.43\pm 0.23$ $81.98\pm 0.48$ $69.29\pm 0.57$ $59.36\pm 2.43$ $46.93\pm 2.21$ $39.68\pm 1.15$
Entropy $94.13\pm 0.30$ $92.39\pm 0.40$ $75.34\pm 4.32$ $60.66\pm 6.23$ $46.17\pm 5.96$ $35.85\pm 4.33$ $29.28\pm 4.29$
Forgetting $94.76\pm 0.38$ $94.34\pm 0.17$ $57.72\pm 1.17$ $35.21\pm 0.87$ $30.87\pm 0.57$ $27.22\pm 0.49$ $23.62\pm 0.74$
EL2N $94.75\pm 0.32$ $94.10\pm 0.45$ $46.41\pm 4.84$ $22.40\pm 0.62$ $15.66\pm 0.18$ $13.05\pm 0.20$ $12.75\pm 0.56$
AUM $\underline{94.90\pm 0.17}$ $\underline{94.38\pm 0.20}$ $51.98\pm 1.32$ $30.56\pm 0.42$ $22.92\pm 0.38$ $18.10\pm 0.47$ $14.50\pm 1.35$
Variance $90.45\pm 0.09$ $85.81\pm 0.39$ $64.90\pm 0.53$ $53.64\pm 1.02$ $44.45\pm 0.77$ $36.61\pm 0.70$ $31.46\pm 0.92$
$\bm{k}$-means $93.87\pm 0.33$ $93.19\pm 0.36$ $83.68\pm 0.94$ $72.46\pm 0.51$ $62.68\pm 1.12$ $48.35\pm 2.29$ $38.73\pm 0.87$
$\bm{k}$-DPP $93.84\pm 0.10$ $92.72\pm 0.18$ $84.09\pm 1.21$ $74.62\pm 0.40$ $61.95\pm 4.87$ $46.58\pm 1.76$ $35.16\pm 2.92$
SES (Ours) $\mathbf{95.01\pm 0.08}$ $\mathbf{94.50\pm 0.05}$ $\mathbf{88.31\pm 0.13}$ $\mathbf{80.24\pm 0.72}$ $\mathbf{69.82\pm 0.84}$ $\mathbf{54.78\pm 1.36}$ $\mathbf{45.25\pm 0.81}$
@@
Table 11: Results on image classification (CIFAR100). The best one is bold, and the runner-up is underlined.
Dataset CIFAR100 (100%:77.90)
Sampling rate 70% 50% 20% 10% 5% 2% 1%
Random $74.72\pm 0.16$ $70.51\pm 0.43$ $52.29\pm 2.32$ $37.38\pm 0.80$ $23.03\pm 0.48$ $13.38\pm 0.36$ $8.57\pm 0.65$
Moderate $74.61\pm 0.10$ $69.95\pm 0.32$ $49.37\pm 0.58$ $30.53\pm 1.44$ $17.68\pm 0.50$ $9.28\pm 0.34$ $5.97\pm 0.35$
CCS $76.10\pm 0.29$ $73.12\pm 0.21$ $58.39\pm 1.04$ $\underline{44.92\pm 1.81}$ $\underline{28.68\pm 0.99}$ $\underline{15.93\pm 0.62}$ $\underline{10.28\pm 0.54}$
$\mathbb{D}^{2}$ Pruning $76.00\pm 0.56$ $\underline{74.45\pm 0.36}$ $\underline{59.08\pm 0.53}$ $44.78\pm 0.78$ $26.68\pm 0.85$ $13.90\pm 0.66$ $9.61\pm 0.44$
GraphCut $76.64\pm 0.13$ $68.79\pm 0.35$ $25.08\pm 1.10$ $16.46\pm 0.35$ $10.78\pm 0.52$ $7.84\pm 0.35$ $5.94\pm 0.16$
Entropy $76.81\pm 0.12$ $62.52\pm 0.26$ $33.28\pm 7.01$ $21.85\pm 7.11$ $13.90\pm 4.75$ $8.31\pm 2.69$ $5.30\pm 1.71$
Forgetting $76.64\pm 0.13$ $68.75\pm 0.34$ $25.17\pm 1.12$ $16.45\pm 0.37$ $10.70\pm 0.49$ $7.82\pm 0.37$ $5.90\pm 0.09$
EL2N $76.15\pm 0.50$ $66.19\pm 0.48$ $14.85\pm 0.43$ $7.50\pm 0.27$ $5.28\pm 0.12$ $3.83\pm 0.07$ $3.17\pm 0.14$
AUM $\underline{76.84\pm 0.40}$ $67.81\pm 0.59$ $16.66\pm 0.90$ $8.54\pm 0.10$ $5.60\pm 0.18$ $4.35\pm 0.17$ $3.56\pm 0.10$
Variance $73.69\pm 0.14$ $68.42\pm 0.16$ $49.60\pm 0.28$ $34.63\pm 0.42$ $22.70\pm 0.35$ $13.71\pm 0.38$ $9.18\pm 0.52$
$\bm{k}$-means $74.93\pm 0.12$ $72.86\pm 0.17$ $56.86\pm 0.88$ $42.69\pm 0.52$ $26.39\pm 0.73$ $14.92\pm 0.75$ $8.90\pm 0.31$
$\bm{k}$-DPP $75.34\pm 0.45$ $73.16\pm 0.45$ $56.88\pm 1.01$ $42.09\pm 0.41$ $26.55\pm 0.9$ $14.31\pm 0.33$ $8.23\pm 0.27$
SES (Ours) $\mathbf{77.23\pm 0.10}$ $\mathbf{74.63\pm 0.17}$ $\mathbf{61.52\pm 0.76}$ $\mathbf{48.02\pm 0.98}$ $\mathbf{33.39\pm 0.27}$ $\mathbf{19.68\pm 0.51}$ $\mathbf{14.16\pm 0.48}$
Table 12: Results on image classification (ImageNet-1K). The best one is bold, and the runner-up is underlined.
Dataset ImageNet-1K (100%:73.63)
Sampling rate 70% 50% 20% 10% 5% 2% 1%
Random $71.63\pm 0.12$ $69.26\pm 0.17$ $58.90\pm 0.12$ $47.10\pm 0.26$ $34.04\pm 0.42$ $16.56\pm 0.11$ $5.50\pm 0.28$
Moderate $71.33\pm 0.04$ $68.72\pm 0.07$ $55.23\pm 0.27$ $40.97\pm 0.18$ $25.75\pm 0.16$ $11.33\pm 0.31$ $4.52\pm 0.24$
CCS $70.74\pm 0.09$ $69.23\pm 0.09$ $60.04\pm 0.95$ $50.41\pm 0.11$ $36.92\pm 1.62$ $19.92\pm 1.36$ $9.43\pm 0.31$
$\mathbb{D}^{2}$ Pruning $71.29\pm 0.07$ $70.32\pm 0.05$ $58.91\pm 0.04$ $50.81\pm 0.24$ $37.12\pm 0.09$ $18.97\pm 0.05$ $11.23\pm 0.26$
GraphCut $68.91\pm 0.05$ $68.72\pm 0.07$ $55.28\pm 0.17$ $44.79\pm 0.06$ $33.54\pm 0.12$ $20.07\pm 0.21$ $11.49\pm 0.12$
Entropy $70.93\pm 0.74$ $69.21\pm 0.12$ $54.76\pm 0.16$ $38.46\pm 0.34$ $22.78\pm 0.73$ $7.01\pm 0.24$ $1.95\pm 0.15$
Forgetting $70.57\pm 0.04$ $70.46\pm 0.10$ $60.77\pm 0.02$ $48.73\pm 0.12$ $33.86\pm 0.15$ $15.13\pm 0.23$ $5.66\pm 0.30$
EL2N $71.68\pm 0.04$ $65.98\pm 0.12$ $31.90\pm 3.64$ $12.57\pm 0.11$ $6.50\pm 0.36$ $3.25\pm 0.13$ $1.90\pm 0.09$
AUM $69.94\pm 0.30$ $65.36\pm 0.11$ $21.91\pm 0.13$ $10.50\pm 0.19$ $6.42\pm 0.07$ $3.58\pm 0.07$ $2.24\pm 0.05$
Variance $70.12\pm 0.09$ $66.09\pm 0.01$ $35.15\pm 0.09$ $13.85\pm 0.02$ $7.13\pm 0.05$ $4.72\pm 0.02$ $1.81\pm 0.06$
$\bm{k}$-means $70.33\pm 0.07$ $69.47\pm 0.11$ $59.23\pm 0.22$ $48.12\pm 0.21$ $35.51\pm 0.13$ $18.67\pm 0.08$ $9.65\pm 0.33$
$\bm{k}$-DPP $70.84\pm 0.09$ $69.85\pm 0.14$ $59.92\pm 0.33$ $46.10\pm 0.33$ $34.41\pm 0.07$ $16.33\pm 0.15$ $7.49\pm 0.17$
SES (Ours) $\mathbf{72.80\pm 0.03}$ $\mathbf{71.05\pm 0.09}$ $\mathbf{63.24\pm 0.06}$ $\mathbf{53.59\pm 0.09}$ $\mathbf{41.88\pm 0.13}$ $\mathbf{25.59\pm 0.17}$ $\mathbf{13.43\pm 0.37}$
Table 13: Results on text classification (ANLI). The best one is bold, and the runner-up is underlined.
Dataset ANLI (100%:49.25)
Sampling rate 70% 50% 20% 10% 5% 2% 1%
Random $47.08\pm 0.26$ $45.20\pm 0.50$ $42.13\pm 0.54$ $39.52\pm 0.93$ $38.82\pm 1.06$ $37.50\pm 1.29$ $35.96\pm 1.04$
Moderate $46.84\pm 0.31$ $45.11\pm 0.39$ $41.95\pm 0.33$ $40.16\pm 0.85$ $38.99\pm 0.45$ $35.83\pm 1.39$ $33.91\pm 0.62$
CCS $46.56\pm 0.23$ $45.92\pm 0.70$ $41.67\pm 1.39$ $\underline{41.63\pm 0.97}$ $40.33\pm 0.65$ $37.41\pm 0.05$ $\underline{36.82\pm 0.49}$
$\mathbb{D}^{2}$ Pruning $48.56\pm 1.22$ $47.49\pm 0.17$ $42.77\pm 0.36$ $41.43\pm 0.24$ $\underline{40.34\pm 0.07}$ $\underline{37.92\pm 1.44}$ $36.29\pm 0.59$
GraphCut $46.14\pm 0.67$ $44.53\pm 0.55$ $42.12\pm 0.66$ $39.86\pm 0.27$ $38.15\pm 0.74$ $35.44\pm 0.80$ $34.02\pm 0.61$
Entropy $46.32\pm 1.11$ $45.53\pm 0.44$ $41.45\pm 0.33$ $39.67\pm 0.28$ $38.54\pm 0.66$ $36.69\pm 0.53$ $36.40\pm 0.19$
Forgetting $\underline{48.73\pm 0.35}$ $42.29\pm 0.17$ $39.82\pm 0.16$ $38.37\pm 0.60$ $35.95\pm 0.33$ $35.78\pm 0.42$ $35.03\pm 0.77$
EL2N $48.70\pm 0.65$ $47.85\pm 0.98$ $43.14\pm 1.68$ $39.63\pm 2.20$ $37.52\pm 1.05$ $34.33\pm 1.15$ $34.27\pm 0.42$
AUM $47.86\pm 0.27$ $47.58\pm 0.37$ $\underline{43.57\pm 0.71}$ $40.02\pm 1.51$ $34.66\pm 0.73$ $34.16\pm 0.12$ $33.62\pm 0.12$
Variance $47.97\pm 0.30$ $\underline{47.87\pm 0.56}$ $40.70\pm 0.16$ $38.75\pm 0.62$ $33.52\pm 0.03$ $33.50\pm 0.03$ $33.17\pm 0.12$
$\bm{k}$-means $46.48\pm 0.24$ $46.52\pm 1.06$ $42.42\pm 0.72$ $40.57\pm 0.65$ $39.89\pm 0.10$ $36.74\pm 0.29$ $36.11\pm 0.40$
$\bm{k}$-DPP $47.74\pm 0.07$ $47.02\pm 0.02$ $43.44\pm 0.59$ $40.98\pm 0.46$ $40.12\pm 0.26$ $37.44\pm 0.11$ $36.66\pm 0.23$
SES (Ours) $\mathbf{49.00\pm 0.08}$ $\mathbf{48.22\pm 0.54}$ $\mathbf{45.94\pm 0.27}$ $\mathbf{43.63\pm 1.00}$ $\mathbf{41.82\pm 0.43}$ $\mathbf{39.88\pm 0.7}$ $\mathbf{38.16\pm 0.38}$
Table 14: Results on text classification (IMDB). The best one is bold, and the runner-up is underlined.
Dataset IMDB Review(100%:95.90)
Sampling rate 70% 50% 20% 10% 5% 2% 1%
Random $95.25\pm 0.21$ $95.04\pm 0.31$ $93.53\pm 0.68$ $91.89\pm 0.17$ $89.60\pm 1.89$ $83.25\pm 4.96$ $73.31\pm 1.94$
Moderate $95.39\pm 0.19$ $95.31\pm 0.07$ $\underline{94.03\pm 0.73}$ $92.55\pm 0.47$ $90.34\pm 0.43$ $58.34\pm 4.29$ $50.81\pm 0.55$
CCS $95.22\pm 0.18$ $95.35\pm 0.19$ $93.44\pm 0.61$ $91.87\pm 0.23$ $89.89\pm 1.47$ $85.73\pm 4.76$ $82.05\pm 4.21$
$\mathbb{D}^{2}$ Pruning $95.43\pm 0.16$ $\underline{95.40\pm 0.20}$ $93.75\pm 0.46$ $92.44\pm 0.60$ $90.77\pm 0.33$ $\underline{87.40\pm 0.78}$ $80.06\pm 0.85$
GraphCut $95.39\pm 0.15$ $95.21\pm 0.29$ $93.37\pm 0.57$ $91.69\pm 0.66$ $\underline{90.93\pm 0.31}$ $86.85\pm 1.81$ $\underline{82.26\pm 0.90}$
Entropy $95.39\pm 0.31$ $95.17\pm 0.22$ $93.92\pm 0.50$ $\underline{92.57\pm 0.97}$ $90.30\pm 0.63$ $77.16\pm 4.52$ $59.41\pm 4.70$
Forgetting $95.41\pm 0.11$ $95.31\pm 0.17$ $93.66\pm 0.30$ $89.41\pm 0.58$ $58.78\pm 2.28$ $55.81\pm 4.65$ $52.38\pm 1.03$
EL2N $95.29\pm 0.19$ $95.34\pm 0.21$ $91.31\pm 0.19$ $60.29\pm 0.09$ $49.88\pm 1.35$ $47.18\pm 3.30$ $43.74\pm 0.81$
AUM $95.27\pm 0.66$ $95.23\pm 0.16$ $90.60\pm 0.13$ $55.68\pm 0.21$ $49.81\pm 1.50$ $43.13\pm 4.53$ $36.29\pm 1.45$
Variance $\underline{95.44\pm 0.16}$ $\underline{95.40\pm 0.20}$ $93.75\pm 0.46$ $92.44\pm 0.60$ $90.77\pm 0.33$ $87.40\pm 0.78$ $80.06\pm 0.85$
$\bm{k}$-means $95.34\pm 0.17$ $95.22\pm 0.23$ $93.98\pm 0.54$ $92.49\pm 0.74$ $90.21\pm 0.88$ $86.96\pm 1.84$ $81.51\pm 6.72$
$\bm{k}$-DPP $95.29\pm 0.37$ $95.13\pm 0.21$ $93.76\pm 0.59$ $92.50\pm 0.58$ $89.97\pm 1.42$ $86.14\pm 2.53$ $77.07\pm 3.31$
SES (Ours) $\mathbf{95.60\pm 0.16}$ $\mathbf{95.40\pm 0.12}$ $\mathbf{94.57\pm 0.09}$ $\mathbf{92.96\pm 0.56}$ $\mathbf{91.42\pm 0.24}$ $\mathbf{88.58\pm 0.47}$ $\mathbf{83.14\pm 2.05}$
@@
Random $74.02\pm 0.26$ $72.10\pm 0.32$ $65.45\pm 0.45$ $57.56\pm 0.45$ $43.47\pm 0.67$ $18.78\pm 0.21$ $9.24\pm 0.96$
Moderate $73.42\pm 0.46$ $72.03\pm 0.52$ $65.12\pm 0.57$ $54.71\pm 0.10$ $40.20\pm 0.54$ $15.97\pm 0.36$ $5.13\pm 0.49$
CCS $\underline{74.64\pm 0.15}$ $72.27\pm 0.23$ $65.72\pm 0.59$ $57.35\pm 0.63$ $39.01\pm 0.81$ $17.26\pm 0.62$ $8.49\pm 0.57$
$\mathbb{D}^{2}$ Pruning $74.46\pm 0.52$ $\underline{72.55\pm 0.23}$ $65.59\pm 0.84$ $55.73\pm 0.34$ $44.04\pm 0.41$ $19.16\pm 0.14$ $10.75\pm 0.75$
GraphCut $67.45\pm 0.35$ $64.15\pm 0.54$ $53.12\pm 0.30$ $38.29\pm 0.12$ $26.81\pm 0.32$ $8.56\pm 0.64$ $8.16\pm 0.54$
AL-MDN $74.51\pm 0.35$ $70.36\pm 0.17$ $65.26\pm 0.82$ $54.51\pm 0.94$ $30.85\pm 1.71$ $12.33\pm 1.44$ $8.97\pm 1.33$
$\bm{k}$-means $74.22\pm 0.31$ $72.35\pm 0.19$ $65.52\pm 0.98$ $57.01\pm 0.49$ $43.35\pm 0.60$ $16.96\pm 0.78$ $5.16\pm 0.69$
$\bm{k}$-DPP $74.19\pm 0.38$ $71.99\pm 0.27$ $65.39\pm 0.53$ $56.80\pm 0.17$ $43.92\pm 0.53$ $18.20\pm 0.44$ $10.50\pm 0.62$
SES (Ours) $\mathbf{75.20\pm 0.29}$ $\mathbf{73.33\pm 0.33}$ $\mathbf{66.52\pm 0.41}$ $\mathbf{59.52\pm 0.14}$ $\mathbf{45.92\pm 0.46}$ $\mathbf{23.39\pm 0.22}$ $\mathbf{16.15\pm 0.62}$